John is thrice as old as her daughter Alice. Ten years ago John was five times as old as Alice. How old was they ten years ago.
Answer by ankor@dixie-net.com(22740)

John is thrice as old as her daughter Alice.
j = 3a
Ten years ago John was five times as old as Alice.
j - 10 = 5(a - 10)
j - 10 = 5a - 50
j = 5a - 50 + 10
j = 5a - 40
replace j with 3a (from the 1st statement)
3a = 5a - 40
3a - 5a = -40
-2a = -40
a = -40/-2
a = 20 yrs, A's age now
"How old were they ten years ago?"
A was 10 yrs old 10 years ago
:
Find J's age
3(20) = 60 yrs is J's present age
therefore
60 - 10 = 50 yrs 10 yrs ago